Buy-Rite Pharmacy has purchased a small auto for delivering prescriptions. The auto was purchased for $23,000 and will have a 6-year useful life and a $5,200 salvage value. Delivering prescriptions (which the pharmacy has never done before) should increase gross revenues by at least $33,200 per year. The cost of these prescriptions to the pharmacy will be about $27,400 per year. The pharmacy depreciates all assets using the straight-line method. The payback period for the auto is closest to (Ignore income taxes.):

rev: 05_18_2018_QC_CS-127512

Multiple Choice

4 years

3.1 years

4.9 years

4.7 years
Answer #1
Net annual cash flows = Gross revenues-Costs = 33200-27400= $5800
Payback period = Investment cost/Net annual cash flows
Payback period = 23000/5800= 4 years
Option 1 is correct